OpenStudy (emilyjones284):

Rolle's Theorem f(x)=(x-2)(x-3)(x-4) on [2,4]

OpenStudy (perl):

did you satisfy the hypotheses of rolle's theorem, first?

OpenStudy (perl):

it must be differentiable on (2,4) and continuous on [2,4] , and it is since it is a polynomial

OpenStudy (emilyjones284):

thats what I don't know how to do

OpenStudy (emilyjones284):

Oh okay I wasn't really sure that it was a polynomial

OpenStudy (perl):

its a polynomial in factored form (linear factors)

OpenStudy (emilyjones284):

so then plug in 2 and 4

OpenStudy (emilyjones284):

and then take the derivative to find c?

OpenStudy (perl):

first plug in f(2) and f(4)

OpenStudy (perl):

rolles theorem says we must have f(a) = f(b)

OpenStudy (emilyjones284):

okay well I got -2 for f(2) and 0 for f(4) so rolle's theorem does not apply right?

OpenStudy (perl):

should be zero for f(2)

OpenStudy (perl):

if you plug in 2 , you have (2-2) * whatever = 0

OpenStudy (emilyjones284):

oops i see what I did wrong

OpenStudy (emilyjones284):

so then i just take the derivative and find c right?

OpenStudy (perl):

so now since we have f(2) = f(4) , then there exists a value c inside (2,4) such that f ' (c) = 0

OpenStudy (perl):

correct, take derivative

OpenStudy (emilyjones284):

the derivative is just 1 right?

OpenStudy (perl):

the derivative should not be 1

OpenStudy (emilyjones284):

but the derivative for each is 1 isn't it?

OpenStudy (perl):

yes but youre using product rule incorrectly

OpenStudy (emilyjones284):

so then it's just (1)(1)(1)

OpenStudy (emilyjones284):

what am I doing wrong?

OpenStudy (perl):

well

OpenStudy (perl):

you want to use product rule?

OpenStudy (perl):

here is the product rule ( u * v ) ' = u' * v + u * v '

OpenStudy (emilyjones284):

yes I think

OpenStudy (perl):

we have three factors here though

OpenStudy (emilyjones284):

ohhh I thought i couldve just taken the derivative of each

OpenStudy (perl):

( u * v * w ) ' = u ' * v * w + u * v ' * w + u * v * w '

OpenStudy (perl):

i dont think this is the easiest way to differentiate it , though

OpenStudy (emilyjones284):

ohh there's an easier way?

OpenStudy (perl):

yes we can expand the original f(x)

OpenStudy (anonymous):

product rule?!

OpenStudy (perl):

technically you could use the product rule :)

OpenStudy (perl):

but will be more work

OpenStudy (emilyjones284):

so what should I do?

OpenStudy (perl):

f(x)=(x-2)(x-3)(x-4) = ( x^2 -2x -3x + 6) (x-4)

OpenStudy (perl):

so first expand (x-2)(x-3) using foil

OpenStudy (emilyjones284):

oh and then use the product rule from there?

OpenStudy (perl):

then multiply by x - 4

OpenStudy (perl):

(x-2)(x-3)(x-4) = ( x^2 -2x -3x + 6) (x-4) = ( x^2 - 5x + 6) ( x - 4) = x^3 - 5x^2 + 6x - 4x^2 +20x - 20

OpenStudy (perl):

f(x) = x^3 -8x^2 + 26x - 20

OpenStudy (perl):

ok? foil then foil again

OpenStudy (emilyjones284):

and then I take the derivative?

OpenStudy (perl):

yes

OpenStudy (perl):

now its easy to take the derivative

OpenStudy (emilyjones284):

ohh okay thank you so much